March 5, 2009
- Our PCR from Tuesday did not work, it sat in the PCR machine for 14 hours before it was actually ran. So we obtained no good results. However, other groups are having problems with PCR and getting the positive control to work so today we ran positive controls to ensure that our mastermix is working properly. We used two different plasmids (1A7 and MM10) and three different temperature gradients (48, 52, and 56 degrees Celsius). If this works then we can continue on with running PCR with our onion DNA using 5 different gradients with our two primers. Otherwise we will have to use different mastermix or find a solution to why the positive controls won't work.
- Today we also repeated the gel electrophoresis for the plasmids that we had done on Tuesday. We also used a smaller amount of the digest (18 microliters instead of 24 microliters). This time the gel worked, showing that we did have plasmid DNA.